CIS 221 Software Development Using Components CIS 222 Devlopment of Software Components CIS321 Design and Analysis of Algorithms CIS 360 Introduction to Computer Systems CIS 560 Systems Software Design CIS 601 Social and Ethical Issues in Computing CIS 625 Automata and Formal Languages CIS 655 Concepts of Programming Languages CIS 660 Introduction to Operating Systems